Question: Apply the Linear dependence and Linear Independence in vector definitions and show that the given vectors are Linearly dependant or independent vectors in R4.

V1 = (1, 3, -1, 4)T, V2= (3, 8, -5, 7)T, V3= (2, 9, 4, 23)T.


1
Expert's answer
2022-02-03T11:31:19-0500

Solution:

Given, V1 = (1, 3, -1, 4)T, V2= (3, 8, -5, 7)T, V3= (2, 9, 4, 23)T

We have,

a(1, 3, -1, 4)T +b(3, 8, -5, 7)T +c(2, 9, 4, 23)T=(0 0 0 0)T

Or we can write as follows:

"a+3b+2c=0\ ...(i)
\\ 3a+8b+9c=0\ ...(ii)
\\ -a-5b+4c=0\ ...(iii)
\\ 4a+7b+23c=0\ ...(iv)"

On solving (i), (ii) and (iii), we get

"a=-11c,\:b=3c \ ...(v)"

Put (v) in (iv).

"4a+7b+23c=0
\\ \Rightarrow 4(-11c)+7(3c)+23c=0
\\ \Rightarrow -44c+21c+23c=0
\\ \Rightarrow 0=0"

It means it is true for all c.

Thus, we don't have any of a,b,c non-zero.

So, given vectors are linearly dependent.